OpenStudy (dcher03):

Please help, I will medal and Fan best answer: A number to the 12th power divided by the same number to the 9th power equals 125. What is the number?

OpenStudy (will.h):

Omg I'm latterly on the phone and solving this question in my mind so this is a rational equation X^12/x^9 = 125 X= 5

OpenStudy (will.h):

If it doesn't make sense simplify you'll get X^3 = 125 Apply cube root and you'll get x= 5

OpenStudy (yanasidlinskiy):

Write it in this form:\[\frac{ x^{12} }{ x ^{9} } = 125\] X is the number you are trying to find. This can be simplified to: \[x ^{12-9}=125\rightarrow x ^{3}\]

OpenStudy (dcher03):

sorry, my computer's kinda slow, thanks 2 both of u : )

OpenStudy (yanasidlinskiy):

What number cubed is 125? To find this, cube root both sides: \[x=\sqrt[3]{125} = 5\]

OpenStudy (will.h):

I already said that lol.. You welcome.. :)